### v1

# Strong Haken via Thin Position

Abstract.
We use thin position of Heegaard splittings to give a new proof of Haken’s Lemma that a Heegaard surface of a reducible manifold is reducible and of Scharlemann’s “Strong Haken Theorem”, that a Heegaard surface for a 3-manifold may be isotoped to intersect a given collection of essential spheres and discs in a single loop each. This article could also serve as an introduction to the theory of generalized Heegaard surfaces and it includes a careful study of their behaviour under amalgamation.
1. Introduction
A surface $H$ in a closed, connected, orientable 3-manifold $M$ is a Heegaard surface if $M$ can be obtained (up to homeomorphism) by gluing together two handlebodies via a homeomorphism of their boundaries such that $H$ is the image of the boundaries. This definition can be generalized to 3-manifolds with boundary, by allowing $H$ to divide $M$ into two compressionbodies, rather than handlebodies. The Heegaard genus $\operatorname{g}(M)$ of $M$ is the minimal genus of a Heegaard surface in $M$. A 3-manifold $M$ is reducible if there is an embedded sphere $S\subset M$ such that $S$ is not the boundary of a 3-ball in $M$. A reducing sphere for a Heegaard surface is an essential sphere $S\subset M$ such that $S\cap H$ consists of a single curve in $H$. A Heegaard surface is reducible if there exists a reducing sphere.
One of the fundamental results of Heegaard splitting theory is:
Lemma 1.1 (Haken’s Lemma (1968) [Haken]).
A Heegaard surface $H$ for a compact, orientable reducible 3-manifold $M$ is reducible.
From this it follows easily that:
Corollary 1.2 (Heegaard genus is additive).
If $M_{1}$ and $M_{2}$ are compact and orientable, then $\operatorname{g}(M_{1}\#M_{2})=\operatorname{g}(M_{1})+\operatorname{g}(M_{2})$.
Casson and Gordon [CG] adapted Haken’s result so that it also applies to discs:
Lemma 1.3 (Disc version of Haken’s Lemma).
If $H$ is a Heegaard surface for a compact orientable 3-manifold $M$ with boundary such that $M$ contains an essential disc $D$, then there exists a disc $D^{\prime}$ with the same boundary as $D$, such that $D$ and $H$ intersect in a single simple closed curve.
As with Haken’s Lemma, this has the corollary:
Corollary 1.4.
If $M$ is the boundary connected sum of 3-manifolds $M_{1}$, and $M_{2}$, then $\operatorname{g}(M)=\operatorname{g}(M_{1})+\operatorname{g}(M_{2})$. Likewise, if $M$ is obtained by attaching a 1–handle to the the boundary of a compact, connected, orientable $M^{\prime}$, then $\operatorname{g}(M)=\operatorname{g}(M^{\prime})+1$.
Noting that Haken’s result only guarantees existence, in 2020, Scharlemann considered the question of whether if $S$ is a fixed sphere or disc, a Heegaard surface $H$ can necessarily be isotoped to intersect the given sphere or disc $S$ in a single simple closed curve. He showed:
Theorem 1.5 (Strong Haken Theorem [Scharlemann]).
Suppose that $\mathcal{S}$ is the union of pairwise disjoint essential spheres and discs in a compact, connected, orientable 3-manifold $M$. If $H$ is a Heegaard surface for $M$, then $H$ can be isotoped so that $H$ intersects each component of $\mathcal{S}$ in a single simple closed curve.
Scharlemann’s proof is quite involved and proceeds by considering the intersections between a spine for one of the handlebodies in the Heegaard splitting given by $H$ and the surfaces in the collection $\mathcal{S}$. In [ST-Wald] (see also [SSS]), Scharlemann and Thompson use a similar perspective to prove Haken’s Lemma and Waldhausen’s classification of Heegaard splittings of $S^{3}$ [Waldhausen-3sph]. (Another earlier proof of Haken’s Lemma using normal surface theory was given by Johannson [Johannson].) Scharlemann’s approach also allows him to prove a version of the theorem when $\partial M$ contains spheres. That adds some complication, since it is then possible for an essential sphere to be altogether disjoint from a Heegaard surface.
In 2021, Hensel and Schultens [HeSc] gave a much shorter proof of the Strong Haken Theorem for a closed 3-manifold $M$. Their proof uses “sphere complexes,” a construction arising from recent work in geometric group theory. In this paper, we give another, completely different, proof of the Strong Haken Theorem using thin position for 3-manifolds; an idea that goes back to Scharlemann-Thompson [ST]. Although our proof occupies more pages, there is a certain simplicity to the basic idea. Along the way we develop some Heegaard splitting techniques that may be of independent interest. It is also worth noting that results such as the Strong Haken Theorem are by no means obvious. Recently, the author [Taylor] showed that there is no such result if $H$ and $\mathcal{S}$ are required to be invariant under a smooth finite group action. Indeed, equivariant Heegaard genus can be sub-additive, additive, or super-additive under equivariant connect sums of 3-manifolds.
2. Outline
Heegaard splitting splitting theory has reached considerable maturity as a mathematical subject. A key insight that is the basis of many modern results is the realization (perhaps originating in [G3, Kobayashi]) that strongly irreducible Heegaard splittings can isotoped to intersect essential surfaces in circles of intersection that are mutually essential or mutually inessential. This result does not use Haken’s Lemma; we sketch the proof in Lemma 3.6. Before expounding on some of the details, here is a brief sketch of the argument, for the experts,. For simplicity, assume that $H$ is a Heegaard surface for a closed, connected, orientable, reducible 3-manifold $M=M_{1}\#M_{2}$ where both $M_{1}$ and $M_{2}$ are prime and irreducible. Let $S\subset M$ be an essential sphere. In this simplified setting $S$ is unique up to isotopy. (Consequently, Haken’s Lemma will directly imply the Strong Haken Theorem, however to give an argument that applies more generally we won’t use that fact.) Thin $H$ to a generalized Heegaard surface $\mathcal{H}$ consisting of thick surfaces $\mathcal{H}^{+}$ and thin surfaces $\mathcal{H}^{-}$. This can be done so that $\mathcal{H}^{-}$ is incompressible and each component of $\mathcal{H}^{+}$ is strongly irreducible. The incompressibility of $\mathcal{H}^{-}$ and the strong irreducibility of $\mathcal{H}^{+}$ can be used to isotope $\mathcal{H}$ to be disjoint from $S$. It then follows that some component of $\mathcal{H}^{-}$ is an essential sphere $P$. By our simplifications, the surface $S$ must be parallel to $P$ and surgery on $P$ produces $M_{1}$ and $M_{2}$. On each side of $P$ amalgamate $\mathcal{H}$ to Heegaard surfaces $H_{1}$ and $H_{2}$ for $M_{1}$ and $M_{2}$. We may then amalgamate $H_{1}\cup P\cup H_{2}$ to a Heegaard surface $H^{\prime}$. We perform this amalgamation carefully. Start by sliding the feet of the 1-handles, of the compressionbody between $P$ and $H_{1}$ so that when we extend them to lie on $H_{2}$, we need only extend a single foot. This ensures that $P$ (and, in this simplified setting, $S$) intersect $H^{\prime}$ in a single simple closed curve. We may then appeal to a result of Lackenby [Lackenby, Prop. 3.1] which guarantees that $H$ is isotopic to $H^{\prime}$. In the general setting, where $M$ admits multiple essential spheres and discs, and especially when some of those spheres may be nonseparating, we need to do more work to implement the concept.
Here is a bit more information on the overall strategy. A Heegaard surface is weakly reducible if it admits disjoint compressing discs on opposites sides. A Heegaard surface is strongly irreducible if it is not weakly reducible. Thin position for 3-manifolds is technique for acquiring topological information from a weakly reducible Heegaard splitting of a 3-manifold $M$. It involves replacing a Heegaard surface with a (likely disconnected surface) called a generalized Heegaard surface; each component of a generalized Heegaard surface is either a thick surface or a thin surface. The thick surfaces are Heegaard surfaces for the regions of $M$ complementary to the thin surfaces. The basic idea (originating with Casson and Gordon [CG]) is to decompose a weakly reducible Heegaard surface into three mutually disjoint surfaces (each is potentially disconnected), two of which are thick and one of which is thin. Thinking of a Heegaard splitting as a sequence of handle attachments, where 0-handles and 1-handles are attached before 2-handles and 3-handles, this operation (called “weak reduction”) is equivalent to attaching a 2-handle (and possibly 3-handle) prior to attaching a 1-handle (and possibly a 0-handle). If one or both of the thick surfaces are themselves weakly reducible we can repeat the operation. The end result is a collection of thick and thin surfaces. If it is ever the case that a thick and thin surface are parallel, we remove them from the collection. This way of creating and removing thick and thin surfaces is called thinning. Scharlemann and Thompson [ST] produce a complexity that decreases under thinning; a collection of thin and thick surfaces is thin if it minimizes their complexity. In a thin collection, the thin surfaces are incompressible and the thick surfaces are strongly irreducible. (In [ST, Rule 5], the authors appeal to Lemma 1.3 for this fact; we replace it with Lemma 3.6.) Scharlemann and Thompson observe that their complexity is (in a certain sense) additive under connected sum. Although they do not indicate what they had in mind for a proof, it is likely they intended to use Haken’s Lemma, together with the prime decomposition theorem for 3-manifolds. Our route on the other hand, proves Haken’s Lemma and the Strong Haken Theorem using (in a certain sense) additivity.
We prove Haken’s Lemma in Theorem 5.12 using thin position machinery and its inverse, amalgamation, a technique due to Schultens [Schultens]. Amalgamation converts a generalized Heegaard surface back into a Heegaard surface. In principle, there are innumerable different ways of amalgamating a generalized Heegaard surface. However, the aforementioned result of Lackenby will show that the results are all isotopic. This allows us, in Section 5, to make certain choices during the amalgamating process to control the intersection with thin spheres.
Our proof of the Strong Haken Theorem requires one additional technique, we call juggling: it allows thin spheres to be moved to different places with respect to the other components of a generalized Heegaard surface. We explain it in Section 6.
The proof of the Strong Haken Theorem is then carried out at the end of Section 6 and in Section 7 as follows. We are given a Heegaard surface $H$ for a 3-manifold that contains essential spheres and discs. We are also given a mutually disjoint collection $\mathcal{S}$ of essential spheres and discs. We thin $H$ to arrive at a thin generalized Heegaard surface $\mathcal{H}$. We use juggling to ensure it is disjoint from $\mathcal{S}$. We then carefully amalgamate $\mathcal{H}$ to a Heegaard surface $J$ and appeal to Lackenby’s result to know that $H$ is isotopic to $J$.
2.1. A word on notation
Unless otherwise specified all 3-manifolds $M$ are compact, connected, and orientable. All surfaces are compact and orientable. In a directed graph (aka digraph) a coherent cycle (resp. coherent path) is an embedded oriented loop (resp. path) such that the orientation of the loop (resp. path) coincides with the orientation of each edge it traverses. As in graph theory, a digraph is acyclic if it does not contain a coherent cycle; the underlying undirected graph of an acyclic digraph may contain cycles. We let $I=[0,1]$. If $D$ is the union of pairwise disjoint discs properly embedded in $M$, we let $M|_{D}$ denote the result of $\partial$-reducing $M$ along $D$. The scars from $D$ are the discs in $\partial M|_{D}$ that are the feet of the 1-handles with cocores $D$ removed from $M$. Similarly, if $S\subset M$ is a sphere, we let $M|_{S}$ denote the result of surgery along $S$, that is the result of removing a regular neighborhood of $S$ from $M$ and then gluing in 3-balls to the resulting boundary components. In $M\setminus S$, the scars from $S$ are the feet of the copies of $S^{2}\times I$ removed from $M$. In $M|_{S}$, the scars from $S$ are the 3-balls that were glued to the scars in $M\setminus S$.
3. Compressionbodies and Generalized Heegaard Surfaces.
A compressing disc for a surface $S$ is an embedded disc $D\subset M$ such that:
  1. (1)
    the simple closed curve $\partial D$ is contained in $S$,
  2. (2)
    the interior of $D$ is disjoint from $S$, and
  3. (3)
    the curve $\partial D$ does not bound a disc in $S$.
An s-compressing disc is a generalization of “compressing disc”: the definition is identical except we replace (3) with:
  1. (3’)
    There is no disc $E\subset S$ with $\partial D=\partial E$ such that $D\cup E$ bounds a ball in $M$ with interior disjoint from $S$.
A semi-compressing disc for $S$ is an s-compressing disc that is not a compressing disc. It necessarily has inessential boundary in $S$. If $\Delta$ is the union of mutually disjoint s-discs for $S$, we can compress $S$ to a (possibly disconnected) surface $S|_{\Delta}$ by letting $S|_{\Delta}$ be the result of removing a regular neighborhood $N(\Delta)\approx(\Delta\times I)$ of $\partial\Delta$ from $S$ and then taking the union with $\Delta\times\partial I$. Observe that $S|_{\Delta}$ inherits a normal orientation from $S$ and that each disc in $\Delta$ leaves two scars on $S|_{\Delta}$.
A surface $S$ in a 3-manifold $M$ is:
  • compressible if there is a compressing disc for $S$;
  • weakly reducible if there exist disjoint s-discs $D_{1}$ and $D_{2}$ for $S$, with boundaries on the same component of $S$, such that near $S$, the discs $D_{1}$ and $D_{2}$ lie on opposite sides of the component of $S$ containing their boundaries and at least one of $D_{1}$ or $D_{2}$ is a compressing disc.
  • strongly irreducible if it is not weakly reducible.
A ball compressionbody $C$ is an oriented 3-manifold homeomorphic to a 3-ball. In this case, we let $\partial_{+}C=\partial C$ and $\partial_{-}C=\varnothing$. A product compressionbody $C$ is an oriented 3-manifold homeomorphic (via an orientation-preserving homeomorphism) to $F\times I$ where $F$ is a closed, connected surface. We let $\partial_{+}C$ be the component of $\partial C$ taken to $F\times\{1\}$ and $\partial_{-}C$ be the component taken to $F\times\{0\}$. A trivial compressionbody is either a ball compressionbody or a product compressionbody. A compressionbody is a compact, connected, oriented 3-manifold $C$ with nonempty boundary, such that:
  1. (1)
    One component of $\partial C$ is designated as $\partial_{+}C$ and the union of the other components is designated as $\partial_{-}C$,
  2. (2)
    There is a (possibly empty) collection $\Delta$ of mutually disjoint s-discs for $\partial_{+}C$, such that $C|_{\Delta}$ is the union of trivial compressionbodies and with the designations of $\partial_{+}$ and $\partial_{-}$ also inherited from $C$. (Such a collection $\Delta$ is a complete collection of s-discs for $C$.
A punctured trivial compressionbody is the result of removing the open neighborhood of a finite (possibly empty) collection of points from the interior of a trivial compressionbody. The positive boundary of a punctured product compressionbody is the same as the positive boundary prior to removing the open balls. Observe that a punctured product compressionbody is a compressionbody. If a compressionbody does not admit any compressing discs for its positive boundary it is a punctured trivial compressionbody.
A spine $\Gamma$ for $C$ is the union of $\partial_{-}C$ with a properly embedded 1-dimensional CW-complex in $C\setminus\partial_{+}C$ such that $C$ deformation retracts to $\Gamma$. A spine can be constructed by attaching to $\partial_{-}C$ or to a vertex (if $\partial_{-}C=\varnothing$) the cocores of the 2-handles that form a regular neighborhood of any complete collection of s-discs.
Lemma 3.1.
The following are true for a compressionbody $C$:
  1. (1)
    $\partial_{-}C$ is incompressible in $C$.
  2. (2)
    If $D\subset C$ is the union of mutually disjoint s-discs for $\partial_{+}C$, then there exists a complete collection $\Delta$ of s-discs for $C$ such that $D\subset\Delta$.
  3. (3)
    If $D\subset C$ is the union of mutually disjoint s-discs for $\partial_{+}C$, then $C|_{D}$ is the disjoint union of compressionbodies, with orientations and $\partial_{+}$ and $\partial_{-}$ designations inherited from $C$.
  4. (4)
    $C$ does not contain an s-disc for $\partial_{+}C$ if and only if it is a ball compressionbody or product compressionbody. If $C$ contains an s-disc for $\partial_{+}C$ but does not contain a compressing disc for $\partial_{+}C$, then it is a punctured product compressionbody.
  5. (5)
    If $S\subset C$ is a closed connected incompressible surface in the interior of $C$, then $S$ is separating and separates $C$ into two compressionbodies. One of the compressionbodies has (a copy of) $S$ as its positive boundary and components of $\partial_{-}C$ as negative boundary and the other has $\partial_{+}C$ as its positive boundary and (a copy of) $S$ and components of $\partial_{-}C$ as its negative boundary. The compressionbody containing $S$ as its positive boundary has at most one negative boundary component that is not a sphere; it contains such a component $F$ if and only if $S$ is not a sphere. If $S$ is not a sphere that compressionbody is a punctured product compressionbody between $S$ and $F$.
Proof.
Properties (1) - (3) are well-known and easily proved. See [TT1, Lemma 3.5] for a generalization to the situation when $C$ contains a certain type of properly-embedded graph. The case when we are reducing along a semi-compressing disc requires the use of the lightbulb trick (see [HW] for a description). Property (4) follows from the definition of compressionbody.
For the proof of (5), suppose that $S\subset C$ is a closed, connected, incompressible surface in the interior of $C$. Let $\Delta$ be a complete collection of s-discs for $C$ chosen to intersect $S$ minimally. An innermost disc argument then shows that $S$ intersects only semi-compressing discs in $\Delta$. Let $\Delta^{\prime}\subset\Delta$ be the union of the compressing discs and let $W\subset C|_{\Delta^{\prime}}$ be the component containing $S$. Let $\widehat{W}$ be the result of capping off spherical components of $\partial_{-}W$ with 3-balls. Note that $\widehat{W}$ is either a 3-ball or a product compressionbody. This does not change the incompressibility of $S$. Hence, if $\widehat{W}$ is a 3-ball, then $S$ is a 2-sphere bounding a 3-ball in $\widehat{W}$, and the result follows easily. Suppose that $\widehat{W}$ is a product compressionbody. Waldhausen’s classification of surfaces in product manifolds [Waldhausen-sufflarge] shows that either $S$ is parallel in $\widehat{W}$ to the unique component of $\partial_{-}\widehat{W}$ or is a sphere bounding a 3-ball in $\widehat{W}$. Conclusion (5) follows by reconstituting $C$ from $\widehat{W}$. ∎
Remark 3.2.
As in Figure 1, $\partial_{-}C$ may admit a semi-compressing disc. This observation will necessitate some extra work in what follows. However, Lemma 3.3 shows that such discs still cut off compressionbodies.
[…]
Figure 1. It is possible for the negative boundary of a compressionbody $C$ to admit a semi-compressing disc $D$.
Lemma 3.3.
Suppose that $C$ is a compressionbody and that $D$ is a semi-compressing disc for $\partial_{-}C$. Then, together with a disc $E\subset\partial_{-}C$, $D$ bounds a compressionbody in $C$ disjoint from $\partial_{+}C$. This compressionbody is the result of removing at least one open 3-ball from a 3-ball with boundary $D\cup E$.
Proof.
By definition of semi-compressing disc, $\partial D$ bounds at least one disc in $\partial_{-}C$. As every sphere in a compressionbody is separating and since $\partial_{-}C$ is incompressible in $C$, there is such a disc $E$ such that the sphere $D\cup E$ bounds a submanifold $W\subset C$ disjoint from $\partial_{-}C$. An innermost disc argument ensures that there is a maximal pairwise disjoint collection of compressing discs for $\partial_{+}C$ disjoint from $D$. Thus, after capping off spherical components of $\partial_{-}C$ (other than the one containing $\partial D$ if it is a sphere) we see that $W$ becomes a 3-ball. Thus, reconstructing $C$, we see that $D\cup E$ bounds a compressionbody in $C$ that is the result of removing at least one open 3-ball from a 3-ball. ∎
The next observation concerning handles is key to the process of amalgamating a generalized Heegaard surface so that it intersects a sphere exactly once. See Figure 2 for an example.
Lemma 3.4 (Monopod Lemma).
Given any compressionbody $C$ with $\partial_{-}C\neq\varnothing$, there exists a collection of s-discs $D\subset C$ such that each component of $C|_{D}$ is either a handlebody or a product compressionbody and each component contains exactly one scar from $D$. Furthermore, if $A\subset C$ is the union of pairwise disjoint vertical annuli and compressing discs for $\partial_{+}C$, we may choose $D$ to be disjoint from $A$.
Proof.
The compressionbody $C$ is obtained from $\partial_{-}C\times I$ by attaching 1-handles to $\partial_{-}C\times\{1\}$. We may slide the ends of the 1-handles onto other 1-handles to ensure that each component of $\partial_{-}C\times\{1\}$ is incident to exactly one foot of the 1-handles. Given $A$, we may start with a choice of 1-handles whose co-cores are disjoint from the annuli components of $A$ and parallel to the disc components of $A$. The slides may then be performed so that the resulting discs $D$ are disjoint from $A$. ∎
Figure 2. By sliding 1-handles, we may find discs separating a compressionbody into product compressionbodies and handlebodies so that each product compressionbody contains a single scar.
A generalized Heegaard surface $\mathcal{H}$ for a 3-manifold $M$ is a closed properly embedded surface such that the following hold:
  1. (1)
    $\mathcal{H}=\mathcal{H}^{+}\cup\mathcal{H}^{-}$, where each of $\mathcal{H}^{+}$ and $\mathcal{H}^{-}$ are unions of components of $\mathcal{H}$ and $\mathcal{H}^{+}\cap\mathcal{H}^{-}=\varnothing$.
  2. (2)
    The closure of the complement of $\mathcal{H}$ in $M$ (denoted $M\setminus\mathcal{H}$) is the union of compressionbodies.
  3. (3)
    $\mathcal{H}^{+}=\bigcup\limits_{C}\partial_{+}C$
  4. (4)
    $\partial M\cup\mathcal{H}^{-}=\bigcup\limits_{C}\partial_{-}C$
where the unions appearing in the last two items are over all the components of $M\setminus\mathcal{H}$ and the equality holds after gluing. A generalized Heegaard surface $\mathcal{H}$ is oriented if each component of $\mathcal{H}\cup\partial M$ is assigned a transverse orientation such that for each compressionbody $C\subset M\setminus\mathcal{H}$, $C$ is a cobordism from $\partial_{-}C$ to $\partial_{+}C$ or vice-versa. We say that a compressionbody $C$ is above (resp. below) $\partial_{+}C$ if the transverse orientation for $\partial_{+}C$ points into (resp. out of) $C$. Two oriented generalized Heegaard surfaces are equivalent if they are ambient isotopic in $M$, via an isotopy preserving orientations. We will only work with oriented generalized Heegaard surfaces in what follows.
Definition 3.5.
Associated to each oriented generalized Heegaard surface $\mathcal{H}$ is a dual digraph. This is a directed graph with a vertex for each component of $M\setminus\mathcal{H}$ and each directed edge from one vertex to another corresponding to an (oriented) component of $\mathcal{H}$. Some edges correspond to thick surfaces and some to thin surfaces. At each vertex, either there is one incoming thick edge and all thin edges are outgoing or there is one outgoing thick edge and all thin edges are incoming. In particular, the requirement that each compressionbody is a cobordism between $\partial_{-}C$ and $\partial_{+}C$ ensures that no edge is a loop. We say that $\mathcal{H}$ is acyclic if its dual digraph is acyclic. (These dual digraphs are essentially the fork complexes of [SSS].) The dual digraph for an acyclic $\mathcal{H}$ may have underlying undirected graph with cycles.
A generalized Heegaard surface is a Heegaard surface if it is connected. In which case, $\mathcal{H}=\mathcal{H}^{+}$ and $\mathcal{H}^{-}=\varnothing$.
Lemma 3.6.
Suppose that $H$ is a strongly irreducible Heegaard surface for $M$ and that $\mathcal{S}$ is a nonempty collection of mutually disjoint essential spheres and discs. Then $H$ can be isotoped so that for each sphere $S_{0}\subset\mathcal{S}$, $H\cap S_{0}$ is the union of loops that are inessential in $H$ and for each disc $S_{0}\subset\mathcal{S}$, $H\cap S_{0}$ is the union of loops, at most one of which is essential in $H$. Furthermore, any component of $H\cap\mathcal{S}$ which is innermost in $\mathcal{S}$ bounds, in $\mathcal{S}$, an s-disc for $H$.
Proof.
This is a by now standard sweepout argument so we simply sketch the proof. For more details, see (for example) [RS]. Let $\Gamma_{-}$ and $\Gamma_{+}$ be spines for the compressionbodies on either side of $M$, transverse to $\mathcal{S}$. Their complement is a product which allows us to define a continuous function $\phi\mskip 0.5mu\colon\thinspace M\to I$. We may perturb it so $\phi|_{\mathcal{C}}$ is Morse. Note that $\phi^{-1}(0)=\Gamma_{-}$, $\phi^{-1}(1)=\Gamma_{+}$ and $H_{t}=\phi^{-1}(t)$ is isotopic to $H$ for any $t\in I\setminus\partial I$. The critical values of $\phi$ divide $I$ into open sub-intervals. Label a subinterval with -- (resp. $+$) if for some (equivalently, any) $t$ in the subinterval has the property that there is an essential loop in $H_{t}\cap\mathcal{S}$ that is essential in $\phi^{-1}(t)$ and which bounds a disc below $H$ (resp. above $H$). As $H$ is strongly irreducible no interval is labelled with opposite signs. Since both $H$ and $\mathcal{C}$ are orientable, it is also the case that no two adjacent intervals have opposite sign.
Consider either the lowest interval or the highest interval. Let $t$ be in the interval and let $C$ be the compressionbody below or above $H_{t}$ respectively. We think of $H_{t}$ as being the frontier of a regular neighborhood of the spine. Let $S_{0}\subset\mathcal{S}$ be a component. Each intersection of $S_{0}$ with $H$ is either a loop in $C$ parallel to $\partial S_{0}$ or a meridian of an edge of the spine. A meridian intersection bounds an s-disc for $H$ in $C$. If some such disc is a compressing disc, the interval is labelled -- (if it is the lowest interval) or $+$ (if it is the highest interval). If not, $H_{t}$ is the surface isotopic to $H$ satisfying the conclusions of the lemma.
Assume, therefore, that both sides of $H$ admit compressing discs, so the lowest interval is labelled -- and the highest is labelled $+$. We conclude that there must be a subinterval with no label. Let $t$ lie in that subinterval. If $H_{t}\cap\mathcal{S}=\varnothing$, we are done, so suppose that $\gamma\subset H_{t}\cap\mathcal{S}$ is a component. As $\mathcal{S}$ is the union of discs and spheres, we may assume that it is innermost in $\mathcal{S}$. Let $D\subset\mathcal{S}$ be the innermost disc it bounds. Since the interval containing $t$ has no label, $\partial D$ is inessential on $H$. If $D$ is not a semi-compressing disc for $H$, we can isotope $H$ so as to remove $\partial D$ from $H\cap\mathcal{H}$. ∎
4. Heegaard surfaces through thick and thin
Following [HS], we reinterpret [ST] in terms of surfaces, rather than handle structures. Suppose that $\mathcal{H}$ is an oriented generalized Heegaard surface and that $H$ is an weakly reducible thick surface. Let $D_{-}$ and $D_{+}$ be unions of pairwise disjoint s-discs for $H$, with $D_{-}$ below $H$ and $D_{+}$ above $H$. Let $H_{\pm}=H|_{D_{\pm}}$ and let $F=H|_{D_{-}\cup D_{+}}$. A small isotopy of $H_{+}$ above $H$ and $H_{-}$ below $H$ makes $H_{-}$, $H_{+}$, and $F$ mutually disjoint. Let $\mathcal{J}^{+}=(\mathcal{H}\setminus H)\cup H_{-}\cup H_{+}$ and let $\mathcal{J}^{-}=\mathcal{H}^{-}\cup F$. Then we say that $\mathcal{J}$ is obtained by an untelescoping $\mathcal{H}$ and that $D_{-}$ and $D_{+}$ are an s-weak reducing pair. It is a weak reducing pair if both $D_{-}$ and $D_{+}$ contain a compressing disc and a (s/2)-weak reducing pair if at least one of $D_{-}$ or $D_{+}$ contain a compressing disc. See Figure 3.
[…]
Figure 3. Untelescoping a thick surface $H$ using a pair of s-discs $D_{-}$ and $D_{+}$.
The following theorem is essentially due to Scharlemann-Thompson, however they do not address the situation with semi-compressing discs; that is explored in depth in [TT1]. In this paper, we somewhat modify both approaches, for simplicity.
Theorem 4.1.
If $\mathcal{H}$ is an oriented generalized Heegaard surface and $\mathcal{J}$ is obtained from $\mathcal{H}$ by untelescoping, then $\mathcal{J}$ is an oriented generalized Heegaard surface. Furthermore, if $\mathcal{H}$ is acyclic, so is $\mathcal{J}$.
Proof.
We leave the proof to the reader; see [ST, HS, TT1] for example. ∎
On the other hand, suppose $\mathcal{H}$ contains a thick surface $H$ and a thin surface $F$ such that $H\cup F$ lie in the boundary of a product compressionbody component $C$ of $M\setminus\mathcal{H}$. We say that $\mathcal{J}=\mathcal{H}\setminus(\mathcal{H}\cup F)$ is obtained by consolidating $\mathcal{H}$.
Lemma 4.2 (Taylor-Tomova [TT1]).
Suppose $\mathcal{H}$ is an oriented generalized Heegaard surface and that $\mathcal{J}$ is obtained from $\mathcal{H}$ by a consolidation. Then $\mathcal{J}$ is an oriented generalized Heegaard surface. Furthermore, if $\mathcal{H}$ is acyclic, so is $\mathcal{J}$.
We identify three thinning moves that may be able to be performed on an oriented generalized Heegaard surface $\mathcal{H}$.
  1. (1)
    consolidation
  2. (2)
    untelescoping using a weak reducing pair
  3. (3)
    untelescoping using an s-weak reducing pair $D_{-}$, $D_{+}$, such that one of $D_{\pm}$ contains a compressing disc and a positive genus component of $H_{\mp}=H|_{D_{\mp}}$ can be consolidated with a component of $\mathcal{H}^{-}$; we also perform that consolidation.
We call (2) a type I untelescoping and (3) a type II untelescoping and consolidation
A thinning sequence is a sequence of consolidations, type I untelescopings, and type II untelescoping and consolidations. If $\mathcal{J}$ is obtained from $\mathcal{H}$ by a (possibly empty) thinning sequence, we say that $\mathcal{H}$ thins to $\mathcal{J}$. If none of the three thinning moves is possible for $\mathcal{H}$, then we say $\mathcal{H}$ is locally thin.
Remark 4.3.
In many papers, including [HS, ST, TT1], additional moves are also allowed, corresponding to the destabilization of Heegaard surfaces. We do not need to make use of those moves.
Theorem 4.4.
Suppose that $\mathcal{H}$ is a generalized Heegaard surface for a 3-manifold $M$. Then every thinning sequence beginning with $\mathcal{H}$ must terminate. Consequently, there exists a locally thin $\mathcal{J}$ such that $\mathcal{H}$ thins to $\mathcal{J}$. If $\mathcal{H}$ is acyclic, so is $\mathcal{J}$.
Proof.
We have already observed in Lemmas 4.1 and 4.2 that if an oriented $\mathcal{H}$ thins to $\mathcal{J}$, then $\mathcal{J}$ is oriented and if $\mathcal{H}$ is acyclic so is $\mathcal{J}$. We need only show that every thinning sequence terminates. Scharlemann and Thompson define the width of a generalized Heegaard splitting. In our context, this is a complexity for a generalized Heegaard surface. Adapting their definition slightly, we let $c(\mathcal{H})$ be the sequence whose entries are $c(H)=(1-\chi(H)/2)$ for each component $H\subset\mathcal{H}^{+}$, arranged in non-increasing order. We note that each entry of the sequence $c(\mathcal{H})$ is a non-negative integer that is 0 if and only if the corresponding surface is a sphere. If $D$ is a collection of pairwise disjoint s-discs for $H$, then each component $H^{\prime}\subset H|_{D}$ has $c(H^{\prime})\leq c(H)$. If $D$ contains at least one compressing disc, then the inequality is strict for every component $H^{\prime}$. If $D$ contains only s-discs, then there is one component $H^{\prime}$ of the same complexity as $H$ and all others are spheres and have complexity zero.
If $\alpha,\beta$ are two such complexity sequences, we say that $\alpha<\beta$ if and only if there exists $k\geq 0$ such that the first $k$ terms of $\alpha$ and $\beta$ coincide and either $\alpha$ does not have a $(k+1)$st term, but $\beta$ does (possibly it is 0) or the $(k+1)$st term of $\beta$ is strictly greater than the $(k+1)$st term of $\alpha$. Observe that complexities are well-ordered.
Consolidation removes a term from a complexity sequence and so decreases it. A type I untelescoping replaces an entry in a complexity sequence with some number (at least two) of other terms, each strictly smaller than the original. Thus, it also decreases complexity. Finally, consider a type II untelescoping and consolidation. Assume the untelescoping using an s-weak reducing pair $D_{-}$, $D_{+}$ such that $D_{-}$ contains only semi-compressing discs and $D_{+}$ contains at least one compressing disc. Let $H\subset\mathcal{H}^{+}$ contain $\partial D_{-}\cup\partial D_{+}$. Since $H$ admits a compressing disc, $c(H)>0$. Then if $H^{\prime}$ is a component of $H_{+}=H|_{D_{+}}$, we have $c(H^{\prime})<c(H)$. If $H^{\prime}$ is a component of $H_{-}=H|_{D_{-}}$, either $c(H^{\prime})=c(H)$ or $c(H^{\prime})=0$. In the former case, we immediately consolidate the component $H^{\prime}$, and so the type II untelescoping and consolidation also strictly decreases complexity. ∎
Lemma 4.5.
Suppose that $\mathcal{H}$ is locally thin oriented generalized Heegaard surface for $M$. Then the following hold:
  1. (1)
    Each component of $\mathcal{H}^{+}$ is strongly irreducible;
  2. (2)
    No component of $M\setminus\mathcal{H}$ is a product compressionbody between components of $\mathcal{H}^{-}$ and $\mathcal{H}^{+}$;
  3. (3)
    If a component $C\subset M\setminus\mathcal{H}$ is a punctured product compressionbody with $\partial_{+}C$ of positive genus and the positive genus component of $\partial_{-}C$ in $\mathcal{H}^{-}$, then the compressionbody on the other side of $\partial_{+}C$ is a punctured product compressionbody;
  4. (4)
    Each component of $\mathcal{H}^{-}$ is incompressible in $M$.
Proof.
Each component of $\mathcal{H}^{+}$ is strongly irreducible in $M\setminus\mathcal{H}^{-}$ as otherwise we could further untelescope $\mathcal{H}$.
Suppose that $C\subset M\setminus\mathcal{H}$ is a punctured product compressionbody with $H=\partial_{+}C$ of positive genus and $F\subset\partial_{-}C\cap\mathcal{H}^{-}$ also of positive genus. $C$ cannot be a product compressionbody for then $\mathcal{H}$ would not be locally thin. Thus $\partial_{-}C$ also contains a sphere and, therefore, a semi-compressing disc. Let $D\subset M\setminus\mathcal{H}$ be the compressionbody on the other side of $H$ from $C$. If $D$ contains a compressing disc for $H$, then it could be isotoped to be disjoint from any semi-compressing disc in $C$. Consequently, we can perform a type II untelescoping and consolidation move. This contradicts the assumption that $\mathcal{H}$ is locally thin. Thus $D$ does not contain a compressing disc for $H$. It must, therefore, be a punctured product compressionbody.
Suppose a component $F$ of $\mathcal{H}^{-}$ is compressible in $M$ via a compressing disc $D$. Choose $F$ and $D$ so that $D$ intersects $F\subset\mathcal{H}^{-}$ minimally in its interior. An innermost disc argument shows that the interior of $D$ is disjoint from $\mathcal{H}^{-}$. Let $W$ be the component of $M\setminus\mathcal{H}^{-}$ containing $D$ and let $H$ be the component of $\mathcal{H}^{+}$ in $W$. By Lemma 3.6, we may isotope $H$ so that $H\cap D$ consists of a single curve that is essential in $H$ and, possibly, a collection of inessential curves. Surgery on $D$ removes those inessential curves, so we may arrange that $D\cap H$ is a single essential curve in $H$. Thus, on one side $C$ of $H$, $D\setminus H$ is a vertical annulus, and on the other side it is a compressing disc for $H$. If $C$ contained a compressing disc for $H$, there would be one disjoint from $D$ and we would contradict the fact that $H$ is strongly irreducible. Thus, $C$ is a punctured product compressionbody. Note that $\partial_{+}C$ has positive genus. By (2), the compressionbody on the opposite side of $H$ is also a punctured compressionbody; however, punctured compressionbodies do not admit compressing discs for their boundaries. Thus, $\mathcal{H}^{-}$ is incompressible. ∎
Remark 4.6.
If $\mathcal{H}$ is locally thin, we cannot guarantee that each component of $\mathcal{H}$ is essential; there may be components that are 2-spheres bounding 3-balls or components that are $\partial$-parallel. Such thin surfaces will not interfere in what follows. As in [ST, HS, TT1], if we allowed certain destabilization operations, we could continue thinning and ensure the thin surfaces were essential.
Corollary 4.7.
Assume that $M$ is a reducible manifold such that no component of $\partial M$ is a sphere. If $\mathcal{H}$ is a locally thin generalized Heegaard surface for $M$, then $\mathcal{H}^{-}$ contains a collection of essential spheres $P$ such that each component of $M|_{P}$ is irreducible. In particular, if $M$ contains a non-separating sphere, then so does $\mathcal{H}^{-}$.
Proof.
Let $P\subset\mathcal{H}^{-}$ be the union of all the spheres in $\mathcal{H}^{-}$ that are essential in $M$. Let $M_{0}$ be a component of $M|_{P}$ and let $\mathcal{H}_{0}$ be those components of $\mathcal{H}$ contained in the interior of $M_{0}$. Using the fact that $\mathcal{H}$ cannot be consolidated and that the existence of spherical components in $\partial_{-}C$ for a compressionbody $C$, does not affect which curves in $\partial_{+}C$ bound compressing discs in $C$, we see that $\mathcal{H}_{0}$ is also thin.
Suppose that $M_{0}$ contains an essential sphere $Q$. Isotope it to intersect $\mathcal{H}_{0}$ minimally. An innermost disc argument shows that there is a component $M^{\prime}_{0}$ of $M_{0}\setminus\mathcal{H}^{-}_{0}$ that either contains $Q$ or contains a semi-compressing disc for a component of $\mathcal{H}^{-}_{0}$. In the latter case, by Lemma 3.3, the compressionbody containing that disc must have a sphere in its negative boundary. That sphere cannot lie in $\partial M_{0}$ so it belongs to $\mathcal{H}^{-}_{0}$. But in that case, it is not essential so it bounds a 3-ball and we can further isotope $Q$ to remove the intersection. Thus, $Q\cap\mathcal{H}^{-}_{0}=\varnothing$. Lemma 3.6 shows that this implies some component of $\mathcal{H}_{0}^{+}$ is weakly reducible, a contradiction.
Since the previous paragraph modifies the sphere $Q$ by surgery, and since every sphere in a compressionbody is separating, if $Q$ were non-separating, we see that $P$ must also have contained a nonseparating sphere. ∎
5. Amalgamation
Suppose that $M$ is a 3-manifold containing an transversally oriented closed surface $F\subset M$ (possibly disconnected) separating $M$ into (possibly disconnected) 3-manifolds $M_{-}$ and $M_{+}$, such that the transverse orientation on $F$ points into each component of $M_{+}$ and out of each component of $M_{-}$ and so that each component of $M_{-}$ and $M_{+}$ is incident to a component of $F$. Suppose that $H_{\pm}$ is the union of oriented Heegaard surfaces for $M_{\pm}$ such that the orientation on $H_{+}$ (resp. $H_{-}$) points away from (resp. towards) $F$. We can amalgamate the surfaces $H_{-}$ and $H_{+}$ across $F$ as follows. See Figure 4.
[…]
Figure 4. The amalgamation of Heegaard surfaces $H_{-}$ and $H_{+}$ across a thin surface $F$.
For notational convenience, Let $U_{\pm},V_{\pm}$ be the union of the compressionbodies on either side of $H_{\pm}$ with $F\subset U_{-}\cap V_{+}$. Choose a complete collection of discs for each of $U_{-}$ and $V_{+}$. When we $\partial$-reduce $U_{-}\cup V_{+}$ along those discs, we obtain product compressionbodies and 3-balls. The union of some of those product compressionbodies is $\Pi=F\times[-1,1]$ where $F\times[-1,0]\subset V_{+}$ and $F\times[0,1]\subset U_{-}$. In both cases, consider $F$ as $F\times\{0\}$. The other product compressionbodies contain $\partial_{-}V_{+}\setminus F$ and $\partial_{-}U_{-}\setminus F$. Let $\Delta_{-}\subset U_{-}$ be those discs in the complete collection that leave scars on $F\times\{1\}$. A small isotopy of the product structure on $\Pi$ ensures that the projection of the scars from $\Delta_{-}$ on $F\times\{1\}$ to $F\times\{-1\}$ completely misses the scars in $F\times\{-1\}$ arising from the complete collection of discs in $V_{+}$. Let $A\subset\Pi$ be the solid tubes corresponding to the projections. Let $X_{-}$ be the union of $V_{-}$ with $A$ and with the components of $U_{-}|_{\Delta_{-}}$ not containing $F$. Let $X_{+}$ be the closure of the complement of $X_{-}$ and let $H=X_{-}\cap X_{+}$. See Figure 4.
Lemma 5.1 (Schultens [Schultens]).
$H$ is a Heegaard surface for $M$.
Proof.
We observe that $X_{-}$ is the result of attaching 1-handles to $\partial_{+}V_{-}$ and so it is a compressionbody. From our construction, we note that the complete set of discs for $U_{-}$ together with a complete set of discs for $V_{-}$ are a complete set of discs for $X_{-}$.
On the other hand, consider a complete set of discs $\Delta_{+}$ for $U_{+}$, with boundaries transverse to $\Delta_{-}$. Then $\partial\Delta_{+}\cap F\times\{1\}$ is a collection of arcs and circles. We may use the product structure on $\Pi$ to extend $\partial\Delta_{+}$ so that it lies on $H$ and is disjoint from the scars on $F\times\{-1\}$ left by the complete collection of discs for $V_{+}$. The union of the complete collection of discs for $V_{+}$ with $\Delta_{+}$ is then a complete collection of discs for $X_{+}$. Consequently, $X_{+}$ is a compressionbody. ∎
Two observations will be useful later:
Lemma 5.2.
Suppose that $M_{-}$ and $M_{+}$ are 3-manifolds such that $F=\partial M_{1}\cap\partial M_{2}$ is nonempty and oriented so that its orientation points into $M_{+}$ and out of $M_{-}$ Let $H_{-}$ and $H_{+}$ be oriented Heegaard surfaces for $M_{-}$ and $M_{+}$ respectively such that the orientation for $H_{+}$ points away from $F$ and the orientation for $H_{-}$ points toward $F$. Let $H$ be the Heegaard surface resulting from amalgamating $H_{+}$ and $H_{-}$ across $F$. Then:
  1. (1)
    If $F$ and $H_{\pm}$ bound a product compressionbody in $M_{\pm}$, then (up to isotopy) $H=H_{\mp}$. That is, the amalgamation is the same as consolidation.
  2. (2)
    If $F$ is a sphere, then $H$ is isotopic to the connected sum of $H_{-}$ and $H_{+}$ as Heegaard surfaces for the 3-manifolds $M_{-}$ and $M_{+}$ obtained by capping off $F$ with 3-balls.
Proof.
We leave the proof to the reader, although see the proof of the Strong Haken Theorem below for an idea that may help with (2). ∎
Definition 5.3.
Suppose that $\mathcal{H}$ is an oriented generalized Heegaard surface and that $H$ is a thick surface. Let $U_{-}$ (resp. $U_{+}$) be the compressionbody of $M\setminus\mathcal{H}$ below (resp. above) $H$. Let $V_{+}$ (resp. $V_{-}$) be the union of some nonempty collection of compressionbodies such that each component of $V_{+}$ (resp. $V_{-}$) shares a negative boundary component with $U_{-}$ (resp. $U_{+}$). Let $J_{\pm}$ be the result of amalgamating $H$ with each component of $\partial_{+}V_{\pm}$. Let $\mathcal{J}$ be the either the result obtained from $\mathcal{H}$ by replacing $H\cup\partial_{+}V_{+}\cup(\partial_{-}V_{+}\cap\partial_{-}U_{-})$ with $J_{+}$ or the result obtained from $\mathcal{H}$ by replacing $H\cup\partial_{+}V_{-}\cup(\partial_{-}V_{-}\cap\partial_{-}U_{+})$ with $J_{-}$. In either case, we say that $\mathcal{J}$ is obtained by an amalgamation of $\mathcal{H}$.
Lemma 5.4.
If $\mathcal{J}$ is obtained from an oriented generalized Heegaard surface $\mathcal{H}$ by an amalgamation, then $\mathcal{J}$ is an oriented generalized Heegaard surface.
Proof.
Use the notation of Definition 5.3. Without loss of generality, consider the case when we amalgamate $H$ with $\partial_{+}V_{+}$. We may also consider just the case when $V_{+}$ consists of a single compressionbody; if it consists of more, just iterate the following argument.
Let $F=\partial_{-}V_{+}\cap\partial_{-}U_{-}$. Let $S=\mathcal{H}^{-}\setminus F$. One component $M_{0}\subset M\setminus S$ contains $F\cup\partial_{+}V_{+}\cup H$ and $F$ separates $M_{0}$ into two submanifolds. One contains $H$ as a Heegaard surface and the other contains $\partial_{+}V_{+}$ as a Heegaard surface. By Lemma 5.1, we may amalgamate $H$ and $\partial_{+}V_{+}$ across $F$ to obtain $J_{+}$, which is a Heegaard surface for $M_{0}$. The surface $J_{+}$ inherits an orientation from $\partial_{+}V_{+}$ that is consistent with the orientation on $H$. Regluing $M_{0}$ to the other components of $M\setminus S$, we see that $\mathcal{J}$ is an oriented generalized Heegaard surface for $M$. ∎
Here is the corollary that is key to our proof of Haken’s Lemma.
Corollary 5.5.
Suppose that $H_{\pm}$ are thick surfaces amalgamated across thin surfaces $F$ to obtain a thick surface $H$. Suppose the amalgamation uses a complete set of discs $\Delta_{-}$ such that after $\partial$-reducing $U_{-}$ using $\Delta_{-}$ each component of $F\times\{1\}\subset U_{-}$ contains a single scar. Then for each component $F_{0}\subset F$, the thick surface $H$ intersects $F_{0}$ in a single simple closed curve.
Proof.
The intersection between $H$ and $F$ occurs on the frontiers of the tubes $A_{-}$. By our choice of $\Delta_{-}$, each component of $F$ intersects a single tube of $A_{-}$ in a single closed curve. ∎
We now turn to the relationship between acyclicity and amalgamation.
Definition 5.6.
A height function on an oriented generalized Heegaard surface $\mathcal{H}$ is a function $f\mskip 0.5mu\colon\thinspace\mathcal{H}\to\mathbb{N}$ which is constant on each component of $\mathcal{H}$ and satisfying:
  • for every component $S\subset\mathcal{H}$, $f(S)$ is odd if and only if $S\subset\mathcal{H}^{+}$;
  • if $C$ is a compressionbody of $M\setminus\mathcal{H}$ with $\partial_{-}C\cap\mathcal{H}^{-}\neq\varnothing$ such that the transverse orientation on $\partial_{+}C$ points out of (resp. into) $C$, then, if $F$ is a component of $\partial_{-}C$ we have $f(F)<f(\partial_{+}C)$ (resp. $f(F)>f(\partial_{+}C)$).
Lemma 5.7.
An oriented generalized Heegaard surface admits a height function if and only if it is acyclic.
Proof.
Suppose that $\mathcal{H}$ is an oriented acyclic generalized Heegaard surface. Each edge of the dual digraph $\Gamma$ corresponds to a component of $\mathcal{H}$. For an edge $S\subset\mathcal{H}$, let $f(S)$ equal the most number of edges in a coherent path from a source of $\Gamma$ to the head of the edge $S$ (including $S$ itself in the count.) Since $\Gamma$ is acyclic and has only finitely many edges, this is well-defined and satisfies the definition of height function.
Conversely, if the dual digraph to $\mathcal{H}$ admits a coherent cycle, it cannot also admit a height function as we would contradict the total ordering of $\mathbb{N}$. ∎
Definition 5.8.
[…]
Lemma 5.9.
If $\mathcal{J}$ is obtained from $\mathcal{H}$ by an amalgamation determined by a height function $f$, then $\mathcal{J}$ is acyclic and $f$ naturally induces a height function for $\mathcal{J}$.
Proof.
Use the notation of Definition 5.8. Without loss of generality assume that $C$ is below $H$. This implies that each $C^{\prime}_{i}$ is above $H_{i}$. Let $C_{i}$ be the compressionbody below $H_{i}$ and $C^{\prime}$ the compressionbody above $H$. Let $H^{\prime}$ be the surface resulting from the amalgamation. Let $U$ and $V$ be the compressionbodies above and below $H^{\prime}$ respectively. Let $\mu=f(H_{i})$ and note this is the same for all $i$.
Define $f^{\prime}\mskip 0.5mu\colon\thinspace\mathcal{J}\to\mathbb{N}$ to be the function which is constant on components of $\mathcal{J}$ and is defined as follows for a component $S\subset\mathcal{J}$:
  • If $S=H^{\prime}$, then $f^{\prime}(S)=\mu$
  • If $S\subset\partial_{-}V\cap\partial_{-}C$, then $f^{\prime}(S)=\mu-1$
  • $f^{\prime}(S)=f(S)$, otherwise.
(If $C$ were above $H$, we would adjust the above by requiring $f^{\prime}(S)=\mu+1$ for any $S\subset\partial_{-}U\cap\partial_{-}C\cap\mathcal{J}$.) See Figure 5 for an example. Observe that $f^{\prime}(S)$ is odd if and only if $S\subset\mathcal{J}^{+}$.
Notice that $\partial_{-}U$ is the union of $\partial_{-}C^{\prime}$ with $\bigcup\limits_{i}\partial_{-}C^{\prime}_{i}\setminus\partial_{-}C$. If $F$ is a component of $\partial_{-}C^{\prime}_{i}\setminus\partial_{-}C$ for some $i$, then
$f^{\prime}(F)=f(F)<f(H_{i})=f^{\prime}(H^{\prime}).$
If $F$ is a component of $\partial_{-}C^{\prime}$, we have
$f^{\prime}(F)=f(F)>f(H)>f(H_{i})=f^{\prime}(H^{\prime}).$
Similarly $\partial_{-}V$ is the union of $\partial_{-}C\setminus\bigcup\limits_{i}\partial_{-}C^{\prime}_{i}$ with $\bigcup\limits_{i}\partial_{-}C_{i}$. If $F$ is a thin surface of $\mathcal{J}$ lying in $\partial_{-}C\setminus\bigcup\limits_{i}\partial_{-}C^{\prime}_{i}$, let $C^{\prime\prime}\subset M\setminus\mathcal{H}$ be the compressionbody on the opposite side of $F$ from $C$. By definition of $\mathcal{A}(C)$,
$f^{\prime}(F)=\mu-1>\mu-2\geq f(H^{\prime\prime})=f^{\prime}(H^{\prime\prime}).$
Finally, if $F$ is a thin surface of $\mathcal{J}$ lying in $\partial_{-}C_{i}$ for some $i$, then
$f^{\prime}(F)=f(F)<f(H_{i})=f^{\prime}(H^{\prime}).$
Thus, $f^{\prime}$ is a height function on $\mathcal{J}$. By Lemma 5.7, $\mathcal{J}$ is acyclic. ∎
[…]
Figure 5. An example of an amalgamation respecting a height function. The numbers indicate the heights of the thick and thin surfaces. The top thick surfaces is amalgamated with the next two thick surfaces on in the left two columns but not the next thick surface in the rightmost column. Those are the two thick surfaces that are amalgamated with the topmost thick surface since they have the largest heights of the thick surfaces immediately adjacent to the topmost thick surface. The fact that the top right thin surface has a higher height is irrelevant.
Corollary 5.10 (cf. Lackenby [Lackenby, Prop. 3.1]).
Suppose that $\mathcal{H}$ is an oriented generalized Heegaard surface with height function $f$. Then there is an oriented Heegaard surface $H$ obtained from $\mathcal{H}$ by a sequence of amalgamations consistent with the induced height functions. Furthermore, up to isotopy, $H$ is independent of both $f$ and the sequence of amalgamations.
Proof.
Let $\mathcal{H}$ be an oriented generalized Heegaard surface with height function $f$ and let $s$ be a sequence of amalgamations consistent with the induced height functions. For any compressionbody $C\subset M\setminus\mathcal{H}$ with $C$ below $\partial_{+}C$, if $F$ is a component of $\partial_{-}C$ with $f(F)<f(\partial_{+}C)$, add $(f(\partial_{+}C)-f(F)+1)/2$ parallel copies of $F$ to $\mathcal{H}$. One component in each copy is a thick surface and one is a thin surface and, together with $F$ and $\partial_{+}C$, the thick and thin surfaces alternate. Observe that we can restore $\mathcal{H}$ by consolidating the parallel copies we just inserted. We can extend both the height function $f$ and the sequence of amalgamations to incorporate these new copies. As amalgamating across a product region is the same as consolidation, this does not affect the result. Consequently, the result of amalgamation consistent with a height function, is independent of the height function, depending only on $\mathcal{H}$ and on the order of thick surfaces chosen in $s$. We now show how to eliminate even this latter dependency.
Let $\mathcal{H}^{\prime}$ be the new generalized Heegaard surface after one of the amalgamations. Its dual digraph is the result of subdividing edges of the dual digraph of $\mathcal{H}$. Let $H_{2i}=f^{-1}(2i)$ and $F_{2i-1}=f^{-1}(2i-1)$ for each $i\in\mathbb{N}$. Observe that if $C$ is a compressionbody below $\partial_{+}(C)$, then
$\partial_{-}C\cap\mathcal{H}^{\prime}\subset f^{-1}(f(\partial_{+}C)-1)$
and if $C$ is above $\partial_{+}C$ then
$\partial_{-}C\cap\mathcal{H}^{\prime}\subset f^{-1}(f(\partial_{+}C)+1).$
We may as well assume, therefore, that for each term of the amalgamation sequence $s$, there is an $i\in\mathbb{N}$ such that the thick surfaces at height $i$ are amalgamated with the thick surfaces at height $i-1$. It then follows from [Lackenby, Prop. 3.1] that $H$ is, up to isotopy, independent of $s$ and of the choice of disc sets used in each amalgamation. ∎
We can now prove additivity of Heegaard genus and Haken’s Lemma. Although the former follows from the latter, it is instructive to prove it independently.
Theorem 5.11 (Additivity of Heegaard genus).
If $M$ contains a nonseparating sphere $S$, then $g(M|_{S})=g(M)-1$ and if a connected 3-manifold $M$ is the connected sum of $M_{1}$ and $M_{2}$, then $g(M)=g(M_{1})+g(M_{2})$.
Proof.
The case when $\partial M$ contains spheres follows relatively easily from the case when it does not; for simplicity, assume that $\partial M$ does not contain spheres. For a generalized Heegaard surface $\mathcal{H}$, let $\operatorname{net}\chi(\mathcal{H})=-\chi(\mathcal{H}^{+})+\chi(\mathcal{H}^{-})$. It is easy to verify that if $\mathcal{H}$ thins to $\mathcal{J}$ or if $\mathcal{J}$ amalgamates to $\mathcal{H}$, then $\operatorname{net}\chi(\mathcal{H})=\operatorname{net}\chi(\mathcal{J})$. We treat the case of separating and non-separating spheres separately; in each case, proving two inequalities.
We start by proving the statement for nonseparating spheres.
Suppose that $M$ contains a non-separating sphere $P$. Let $H$ be a Heegaard surface for $M|_{P}$. One one side of $H$, choose a sphere $F$ bounding a 3-ball $B$. Inside $B$ remove two open 3-balls $B_{1},B_{2}$, creating two spherical boundary components $F_{1},F_{2}$. Let $W=S^{2}\times I$ and let $H^{\prime}\subset W$ be a sphere bounding a 3-ball in $W$. Let $M^{\prime}$ be the result of gluing the components of $\partial W$ to $F_{1}\cup F_{2}$. Then it is possible to assign orientations to $\mathcal{H}=H\cup F_{1}\cup F_{2}\cup H^{\prime}$ so that it is an oriented, acyclic generalized Heegaard surface with $H,H^{\prime}\subset\mathcal{H}^{+}$ and $F_{1},F_{2}\subset\mathcal{H}^{-}$. Notice that $M^{\prime}$ is homeomorphic to $M$ and that
$\operatorname{net}\chi(\mathcal{H})=-\chi(H)-2+4=-\chi(H)+2.$
Amalgamate $\mathcal{H}$ to a Heegaard surface $J$ for $M^{\prime}$. It has genus equal to $g(H)+1$. Thus, $g(M)-1\leq g(M|_{P})$.
Now suppose that $H$ is a Heegaard surface for $M$ of minimal genus. Thin $H$ to a locally thin acyclic oriented generalized Heegaard surface $\mathcal{H}$. By Corollary 4.7, there exists a set of spheres $P\subset\mathcal{H}^{-}$ such that $M|_{P}$ is irreducible and so that $P$ contains a non-separating sphere $P_{0}$. Let $\mathcal{J}=\mathcal{H}\setminus P_{0}$ and observe it is an oriented acyclic generalized Heegaard surface for $M|_{P_{0}}$ with $\operatorname{net}\chi(\mathcal{J})=\operatorname{net}\chi(\mathcal{H})-2=2g(H)-4$. Consistently with a height function, amalgamate $\mathcal{J}$ to a Heegaard surface $J$ for $M|_{P_{0}}$ with $\operatorname{net}\chi(J)=\operatorname{net}\chi(\mathcal{J})$. Consequently, $g(J)=g(H)-1$. Thus, $g(M|_{P_{0}})\leq g(M)-1$. Since, $M|_{P_{0}}$ is homeomorphic to $M|_{P}$, we have $g(M|_{P})=g(M)-1$.
We now prove the additivity result. Assume that $M=M_{1}\#M_{2}$. By the previous result and uniqueness of prime decompositions of 3-manifolds, we may assume that none of $M$, $M_{1}$, $M_{2}$ contain a non-separating sphere.
Let $H_{i}$ be a minimal genus Heegaard surface for $M_{i}$ (with $i=1,2$). In $M$, there is a sphere $F$ separating $H_{1}$ from $H_{2}$ and with $M|_{F}=M_{1}\cup M_{2}$. Give each of $H_{1},H_{2},F$ an orientation so that $H_{1}\cup H_{2}\cup F$ is a generalized Heegaard surface $\mathcal{H}$ with $H_{1},H_{2}\subset\mathcal{H}^{+}$ and $F\subset\mathcal{H}^{-}$. Amalgamating $H_{1}$ and $H_{2}$ across $F$ produces a Heegaard surface $H$ for $M$ with
$22g(H)-2=\operatorname{net}\chi(H)=\operatorname{net}\chi(\mathcal{H})=2g(H_{1})+2g(H_{2})-$ 2 .
Thus, $g(M)\leq g(M_{1})+g(M_{2})$.
Now suppose that $H$ is a Heegaard surface for $M$ of minimal genus. Thin $H$ to a locally thin acyclic oriented $\mathcal{H}$. There exists a set of spheres $P\subset\mathcal{H}^{-}$ such that $M|_{P}$ is irreducible. By uniqueness of prime decompositions of 3-manifolds, we may group the components of $M|_{P}$ into those that are prime factors of $M_{1}$ and those that are prime factors of $M_{2}$. As above, by reconsituting $M_{1}$ and $M_{2}$ from their factors, we may assemble the components of $\mathcal{H}\setminus P$ into oriented, acyclic generalized Heegaard surfaces $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ for $M_{1}$ and $M_{2}$. (This may involve replacing components of $P$ with other thin spheres.) Since every sphere is separating, by the calculation at the start, we see that $g(M)\geq g(M_{1})+g(M_{2})$. Thus, equality holds. ∎
Theorem 5.12 (Haken’s Lemma).
If $H$ is a Heegaard surface for a reducible 3-manifold $M$, then $H$ is reducible.
Proof.
If $\partial M$ contains more than one sphere, it is easy to find an essential sphere intersecting $H$ in a single curve. If $\partial M$ consists of a single sphere, Haken’s Lemma follows from Haken’s Lemma for the manifold obtained by capping off the sphere with a 3-ball. Thus, for simplicity, we may assume that $\partial M$ does not contain a sphere. Let $\mathcal{H}$ be a thin generalized Heegaard surface obtained by thinning $H$ (Theorem 4.4). If $M$ is reducible, then by Corollary 4.7, $\mathcal{H}^{-}$ contains at least one essential sphere $P$. By Lemma 5.1 and Corollary 5.5, we may amalgamate $\mathcal{H}$ to a Heegaard surface $J$ such that $|J\cap P|=1$. By Theorem 5.10, $J$ is isotopic to $H$. That isotopy takes $P$ to a reducing sphere for $H$. ∎
In the next section, we introduce the operation of “juggling” a generalized Heegaard surface. It is an operation that allows us to move spherical thin surfaces from one side of a thick surface to the other side and is used to eliminate inessential curves of intersection between a generalized Heegaard surface and an essential surface. Unfortunately, it can also have the effect of introducing coherent cycles into the dual digraph. Such cycles will always have an edge corresponding to a nonseparating sphere. In order to accommodate those, before proceeding, we generalize our previous work to allow for “spherical self-amalgamation.” See Figure 6 for a depiction.
[…]
Figure 6. An example of spherical self-amalgamation of a Heegaard surface $H$ across a nonseparating sphere $P$. The compressionbody above $H$ in $M\setminus P$ is $A$ and the one below $H$ is $B$. The figure shows the two scars of $P$ in $M|_{P}$, one on either side of $H$. The Heegaard surface $H^{\prime}$ is formed by moving $H$ across one of the scars (by compressing along a semi-compressing disc) and then attaching a tube passing through $P$.
Definition 5.13.
Suppose that $M$ is a 3-manifold containing a nonseparating collection of spheres $P$ and that $H$ is an oriented Heegaard surface for $M\setminus P$ such that for each component $P_{0}\subset P$ the surface $H$ separates the two scars of $P_{0}$. Let $H^{\prime}\subset M$ be the surface constructed as follows. Let $A$ and $B$ be the compressionbodies of $M\setminus P$ above and below $H$ respectively. Let $\Delta_{A}\subset A$ be a collection of semi-compressing discs for $H$ such that $A|_{\Delta_{A}}$ is the disjoint union of $P\times I$ and a compressionbody equivalent to the result of capping off the components of $\partial_{-}A\cap P$ with 3-balls. Assume that $\Delta_{A}$ was chosen so that there is exactly one scar in each component of $P\times I$. Let $\alpha$ be a collection of arcs with endpoints in $H|_{\Delta_{A}}$ and interiors disjoint from $H|_{\Delta_{A}}$ chosen so that:
  1. (1)
    Each arc in $\alpha$ has exactly one endpoint in a scar of $H|_{\Delta_{A}}$.
  2. (2)
    Each scar in $\Delta_{A}$ contains exactly one endpoint of $\alpha$
  3. (3)
    Each component of $\alpha$ is the union of an arc that is a cocore of the 2-handle defined by a disc in $\Delta_{A}$, a vertical arc in one of the components of $P\times I$, and a vertical arc in $A$.
Let $H^{\prime}$ be the result of attaching tubes to $H$ along the arcs of $\alpha$. Note that $H^{\prime}$ is orientable and inherits a transverse orientation from $H$. We say that it is the result of a spherical self-amalgamation of $H$ across $P$.
Lemma 5.14.
If $H^{\prime}$ is obtained from $H$ by a spherical self-amalgamation across $P$, then $H^{\prime}$ is a Heegaard surface for $M$ isotopic to the connected sum of the Heegaard surface $H$ in $M|_{P}$ with $|P|$ copies of the genus 1 Heegaard surface of $S^{1}\times S^{2}$.
Proof.
As above, let $A$ and $B$ be the compressionbodies above and below $H$ in $M\setminus P$, respectively. Let $P\times[-1,1]$ be a regular neighborhood of $P$ in $M$ so that $P\times[-1,0]\subset B$ and $P\times[0,1]\subset A$. The $\partial$-reduction of $A$ along $\Delta_{A}$ results in compressionbodies embedded in $M$. The compressionbody equivalent to the result of capping of the components of $\partial_{-}A\cap P$ does not contain any sphere of $P$. Tubing $H$ along the arcs of $\alpha$ adds 1-handles to the positive boundary of that compressionbody and so it is a compressionbody $A^{\prime}$ embedded in $M$. Hence, $H^{\prime}$ is separating. The other side $B^{\prime}$ of $H^{\prime}$ can be constructed as follows. Begin by letting $B^{\prime}=B$. Let $\Delta_{B}$ be a complete set of compressing and semi-compressing discs for $B$, each disjoint from $\alpha$. The result of $\partial$-reducing $B$ along $\Delta_{B}$ is the union of 3-balls and product compressionbodies whose union contains $P\times[-1,0]$. When we we $\partial$-compress $A$ along $\Delta_{A}$, include $P\times[0,1]$ into $B^{\prime}$. Thus, $P\subset B^{\prime}$ and $B^{\prime}\setminus P$ is equivalent to $B$. The result of $\partial$-reducing $B^{\prime}$ along $\Delta_{B}$ is now the union of 3-balls, product compressionbodies, and components homeomorphic to $(P_{0}\times[-1,1])$, one for each component $P_{0}$ of $P$. Removing a regular neighborhood of $\alpha$ from $B^{\prime}$ to make it part of $A^{\prime}$, converts each of those into a $D^{2}\times[-1,1]$, which is a 3-ball. Undoing the $\partial$-reductions along $\Delta_{B}$, we see that $B^{\prime}$ is a compressionbody.
In $H$, for each arc in $\alpha$ choose an arc in $H$ joining its endpoints. Choose the collection $\beta$ of these arcs to be pairwise disjoint. Note that the regular neighborhood of $P\cup\beta$ consists of $|P|$ copies of $S^{2}\times S^{1}$ with an open 3-ball removed from each. The frontier of a regular neighborhood of $\beta$ witnesses the fact that $H^{\prime}$ is the connected sum of $H$ with $|P|$ copies of the genus 1 Heegaard surface for $S^{2}\times S^{1}$. ∎
Definition 5.15.
Suppose that $\mathcal{H}$ is an oriented generalized Heegaard surface such that each coherent cycle in the dual digraph contains an edge corresponding to a thin sphere. If $P\subset\mathcal{H}^{-}$ is the union of spheres, we say that it is a minimal complete nonseparating collection of spheres if $M\setminus P$ is connected and $\mathcal{H}\setminus P$ in $M\setminus P$ is acyclic. We say that a generalized Heegaard surface $\mathcal{J}$ for $M$ is amalgamation-obtained from $\mathcal{H}$ if there is a minimal complete nonseparating collection of spheres $P\subset\mathcal{H}^{-}$ and a height function $f$ on $\mathcal{H}\setminus P$ for $M\setminus P$ such that $\mathcal{J}$ is obtained by a sequence of amalgamations of $\mathcal{H}\setminus P$ consistent with $f$ followed by spherical self-amalgamation along $P$.
Remark 5.16.
Note that each coherent cycle in the dual digraph to $\mathcal{H}$ contains an edge corresponding to a thin sphere if and only if $\mathcal{H}$ contains a minimal complete nonseparating collection of spheres.
Proposition 5.17.
Suppose that $\mathcal{H}$ is an oriented generalized Heegaard surface containing a minimal complete collection of non-separating spheres. Then there is a Heegaard surface $H$ for $M$ amalgamation-obtained from $\mathcal{H}$. Furthermore, $H$ is unique up to isotopy.
Proof.
The existence of $H$ follows immediately from Lemmas 5.7 and 5.14 and Corollary 5.10. By Corollary 5.10, we need only show that the resulting Heegaard surface does not depend (up to isotopy) on the choice of $P$. Suppose that $P^{\prime}\subset\mathcal{H}^{-}$ is another minimal complete nonseparating collection of spheres. It is enough to consider the case when $P$ and $P^{\prime}$ differ by a single sphere. Let $P_{0}\subset P$ and $P^{\prime}_{0}\subset P^{\prime}$ be the sphere such that $P\setminus P_{0}=P^{\prime}\setminus P^{\prime}_{0}$. Let $M^{\prime}=M\setminus(P\cup P^{\prime}_{0})=M\setminus(P^{\prime}\cup P_{0})$. It has two components, and $\mathcal{H}$ restricts to an acyclic oriented generalized bridge surface in each. Choose a height function for each and amalgamate consistent with the height function to create Heegaard surfaces $H_{1}$ and $H_{2}$. Since $\mathcal{H}$ was oriented, both $H_{1}\cup P_{0}\cup H_{2}$ and $H_{1}\cup P^{\prime}_{0}\cup H_{2}$ are acyclic oriented generalized Heegaard surfaces for $M\setminus P^{\prime}$ and $M\setminus P$ respectively. Amalgamate these to Heegaard surfaces $J^{\prime}$ and $J$ respectively. Observe that each of $J^{\prime}$ and $J$ is isotopic to the result of connect-summing the Heegaard surfaces $H_{1}$ and $H_{2}$. As noted above, performing spherical self-amalmation of $J^{\prime}$ across $P^{\prime}$ is isotopic to the result of performing the connected sum of $J^{\prime}$ with $|P^{\prime}|$ copies of the genus 1 Heegard surface of $S^{2}\times S^{1}$. Similarly, the self-amalgamation of $J$ across $P$ is isotopic to the result of performing the connected sum of $J$ with $|P|$ copies of the genus 1 Heegaard surface of $S^{2}\times S^{1}$. Since $|P|=|P^{\prime}|$, the resulting Heegaard surfaces are isotopic. ∎
6. Juggling
Suppose that $\mathcal{H}$ is a generalized Heegaard surface and that $Q\subset M\setminus\mathcal{H}$ is a sphere. We define an operation, we call a juggle of $Q$. There are three versions; two are depicted in Figure 7.
[…]
Figure 7. The left arrow depicts a juggle (Definition 6.1) where $R$ is a thick surface and $P$ is a thin surface. The right arrow depicts a juggle where $P=R$ is a thick surface. The path $\alpha$ is shown in green for the left juggle. Not shown is the juggle where $P$ is a thick surface and $R$ is a thin surface. The dotted lines simply help delineate the compressionbodies.
Definition 6.1.
Let $Q\subset M\setminus\mathcal{H}$ be a sphere and let $R\subset\mathcal{H}$ be a component. Suppose that $\alpha$ is an embedded path in the interior of $M$ joining $R$ to $Q$, with interior transverse to $\mathcal{H}$, disjoint from $Q$, and intersecting $\mathcal{H}$ exactly once in its interior. Let $P$ be the component of $\mathcal{H}$ intersecting the interior of $\alpha$. (Possibly $P=R$.) Require that one of $P$ or $R$ is a thick surface. Let $Q_{1}$ and $Q_{2}$ be parallel copies of $Q$ in $C$ such that $Q_{2}$ is between $Q_{1}$ and $Q$, each intersecting $\alpha$ exactly once (transversally).
If $P\neq R$, let $P(\alpha)$ be the result of tubing $P$ to $Q_{1}$ along the portion of $\alpha$ from $P$ to $Q_{1}$ and let $R(\alpha)$ be the result of tubing $R$ to $Q_{2}$ along the portion of $\alpha$ from $R$ to $Q_{2}$, so that $R(\alpha)$ is disjoint from $P(\alpha)$. Give $P(\alpha)$ and $R(\alpha)$ the orientations inherited from $P$ and $R$ respectively. If $P=R$, let $P(\alpha)=R(\alpha)$ be the result of first tubing $P$ along the subarc of $\alpha\setminus P$ joining $P$ to $Q_{1}$ and then tubing along the entirety of $\alpha$.
Lemma 6.2.
Suppose that $\mathcal{J}$ is obtained from $\mathcal{H}$ by a juggle. Then the following hold:
  1. (1)
    $\mathcal{J}$ is an oriented generalized Heegaard surface;
  2. (2)
    If $\mathcal{H}$ is acyclic and $\mathcal{H}^{-}$ contains no nonseparating sphere, then $\mathcal{J}$ is acyclic and $\mathcal{J}^{-}$ contains no nonseparating sphere. In which case, any sequence of amalgamations of $\mathcal{H}$ and $\mathcal{J}$ consistent with height functions create isotopic (oriented) Heegaard surfaces for $M$; and
  3. (3)
    If $\mathcal{H}$ is locally thin, then so is $\mathcal{J}$.
Proof.
Use the notation of Definition 6.1. Let $C$ be the compressionbody containing $Q$. Let $C_{Q}\subset C$ be the compressionbody with $\partial_{+}C_{Q}=Q$ that is disjoint from $\partial_{+}C$. Note that $\partial_{-}C_{Q}\subset\partial_{-}C$ is the union of spheres. Note also that $P\subset\partial C$ is not contained in $C_{Q}$. Observe that if we cut $M$ open along $Q$ and paste in a 3-ball to corresponding boundary component of $C$, the arc $\alpha$ guides an isotopy of both $P$ and $R$. Thus, each compressionbody of $M\setminus\mathcal{J}$ is either identical to a compressionbody of $M\setminus\mathcal{H}$, is obtained from $C$ by attaching a 3-ball to $Q$, or is obtained from a compressionbody of $M\setminus\mathcal{H}$ by removing a 3-ball from the interior. It follows that $\mathcal{J}$ is a generalized Heegaard surface. The fact that $\alpha$ intersects exactly $\mathcal{H}$ exactly once in its interior means that the orientation on $\mathcal{H}$ induces an orientation on $\mathcal{J}$.
Consider the effect of the juggle on the dual digraph $\Gamma$ to $\mathcal{H}$. If $P=R$ or if $\partial_{-}C_{Q}\cap\mathcal{H}^{-}=\varnothing$, then the dual digraph is unchanged. Suppose, therefore, that $P\neq R$ and that $\partial_{-}C_{Q}\cap\mathcal{H}^{-}\neq\varnothing$. One of $P,R$ is a thin surface and the other is a thick surface. Let $A$ be the compressionbody containing $P$ be on the opposite side of $P$ from $C$ and let $B$ be the compressionbody containing $R$ on the opposite side of $R$ from $A$. The effect of the juggle is effectively to move $\partial_{-}C_{Q}$ from the compressionbody $C$ to the compressionbody $B$. Thus, any edge of $\Gamma$ dual to a component of $\partial_{-}C_{Q}\cap\mathcal{H}^{-}$ has its end at $C$ slid across the edges dual to $P$ and $R$ to have that end at $B$. If no component of $\partial_{-}C_{Q}\cap\mathcal{H}^{-}$ is nonseparating, each of the corresponding edges of $\Gamma$ separates $\Gamma$ and so $\Gamma$ remains acyclic. Let $F_{1},\ldots,F_{n}$ be the components of $\partial_{-}C_{Q}\cap\mathcal{H}^{-}$. Let $\Gamma_{i}$ be the subgraph of $\Gamma\setminus F_{i}$ not containing the vertex $C$.
Without loss of generality, assume that the orientation on $P$ points into $C$. (If instead it points out of $C$, it is straightforward to adapt the following argument.) We can choose a height function $f$ for $\mathcal{H}$ such that there are $m_{0},m_{1},\ldots,m_{n}\in\mathbb{N}$ with $m_{0}=-1$ such that $f(F_{i})=m_{i}$ and if $S$ is an edge of $\Gamma_{i}$ then $f(S)\in(m_{i-1},m_{i})\subset\mathbb{N}$. Then $f$ is also a height function on $\mathcal{J}$. Amalgamate $\mathcal{H}$ and $\mathcal{J}$ to Heegaard surfaces in such a way that we respect the height function $f$. Do this by amalgamating the thick surfaces in each $\Gamma_{i}$ in the same order and consecutively for $i=1,\ldots,n$. Then amalgamate the thick surfaces not in any $\Gamma_{i}$ in the same order. Finally, perform the remaining amalgamations in the same order. We see that the resulting Heegaard surfaces are isotopic since, in terms of the induced handle structure on $M$, we are performing the same handle attachments in the same order.
Finally, suppose that $\mathcal{H}$ is locally thin. Up to homeomorphism preserving the $\partial_{\pm}$ designations, the difference between the compressionbodies $A$, $B$, $C$, and the compressionbodies $A^{\prime}$, $B^{\prime}$, $C^{\prime}$ is either removing spherical negative boundary components by capping them with 3-balls, or introducing spherical negative boundary components by removing open 3-balls from the interior of the compressionbody. These operations do not affect whether or not a compressionbody is a punctured product compressionbody and they do not affect which curves in a thick surface bound compressing discs. Thus, local thinness is preserved. ∎
We now address the situation when juggling produces coherent cycles in the dual digraph.
Lemma 6.3.
Suppose that $\mathcal{H}$ is an oriented generalized Heegaard surface such that each coherent cycle in its dual digraph contains an edge corresponding to a thin sphere (necessarily nonseparating). Let $\mathcal{J}$ be obtained from $\mathcal{H}$ by a juggle. Then $\mathcal{J}$ is an oriented generalized Heegaard surface such that each coherent cycle in its dual digraph contains an edge corresponding to a thin sphere. Furthermore, if $H$ and $J$ are Heegaard surfaces amalgamation-obtained from $\mathcal{H}$ and $\mathcal{J}$ respectively, then $H$ and $J$ are isotopic.
Proof.
Use the notation of Definition 6.1 and the first two paragraphs of the proof of Lemma 6.2. As before, let $F_{1},\ldots,F_{n}\subset\partial_{-}C_{Q}\cap\mathcal{H}^{-}$, each is a sphere. If none are separating, the result follows immediately from Lemma 6.2, so assume at least one is nonseparating. If the orientation on $P$ points out of $C$, then the edges of the dual digraph $\Gamma$ for $\mathcal{H}$ corresponding to each $F_{i}$ point into the vertex $C$. The effect on the juggle on $\Gamma$ is to take the tips of those edges and slide them in the direction of the edges corresponding to $P$ and $R$. Any coherent cycle in $\Gamma$ after such a slide corresponds to a coherent cycle prior to the slide. Consequently, every coherent cycle for the dual digraph to $\mathcal{J}$ contains an edge corresponding to a thin sphere. This holds even if $P$ or $R$ is a thin sphere, since some $F_{i}$ is nonseparating. Similar to the proof of Lemma 6.2, it is straightforward to see that the Heegaard surfaces amalgamation-obtained from $\mathcal{H}$ and $\mathcal{J}$ are isotopic.
Suppose, therefore, that the orientation for $P$ points into $C$. In this case, the edges of the dual digraph $\Gamma$ for $\mathcal{H}$ corresponding to each $F_{i}$ point out of the vertex $C$. The effect on the juggle on $\Gamma$ is to take the tips of those edges and slide them opposite to the direction of the edges corresponding to $P$ and $R$. This can create cycles in the dual digraph, but each cycle created will contain an edge corresponding to a nonseparating $F_{i}$. Additionally, any coherent cycle in $\Gamma$ prior to the juggle, survives as a coherent cycle after the juggle, possibly with the addition of the edges corresponding to $P$ and $R$.
Let $\mathcal{F}$ be the union of those $F_{i}$ that belong to a coherent cycle in the dual digraph to $\mathcal{J}$ but not in $\mathcal{H}$; however if more than one $F_{i}$ belong to the same cycle include only one from each cycle. Let $P_{H}$ be a minimal complete nonseparating collection of spheres in $\mathcal{H}^{-}$ and expand that to a minimal complete nonseparating collection of spheres in $\mathcal{J}^{-}$ by adding in $\mathcal{F}$. Choose a height function $f$ on $\mathcal{H}\setminus P_{H}$ and let $f^{\prime}$ be its restriction to $\mathcal{H}^{\prime}=\mathcal{H}\setminus(P_{H}\cup\mathcal{F})=\mathcal{J}\setminus(P_{H}\cup\mathcal{F})$. We may amalgamate $\mathcal{H}^{\prime}$ consistent to the height function to a Heegaard surface $L$. The Heegaard surface $H$ is obtained by amalgamating across $\mathcal{F}$ and spherical-self amalgamation across $P_{H}$. The Heegaard surface $J$ is obtained by spherical self-amalgamation across $P_{H}\cup\mathcal{F}$. However, these are simply two ways of forming the connected sum of $L$ with $|P_{H}\cup\mathcal{F}|$ copies of the genus 1 Heegaard surface of $S^{2}\times S^{1}$. Thus, $H$ and $J$ are isotopic. ∎
Corollary 6.4.
Suppose that $H$ is a Heegaard surface for $M$ and that $\mathcal{S}$ is the union of pairwise disjoint essential spheres and discs in $M$. Then there exists an oriented generalized Heegaard surface $\mathcal{H}$ for $M$ such that:
  1. (1)
    Every coherent cycle in the dual digraph to $\mathcal{H}$ contains an edge corresponding to a thin sphere;
  2. (2)
    $H$ is, up to isotopy, amalgamation-obtained from $\mathcal{H}$;
  3. (3)
    Each sphere component of $\mathcal{S}$ is disjoint from $\mathcal{H}$;
  4. (4)
    Each disc component of $\mathcal{S}$ is disjoint from $\mathcal{H}^{-}$ and intersects $\mathcal{H}^{+}$ in a single essential simple closed curve.
Proof.
Let $\mathcal{H}^{\prime}$ be a locally thin, acyclic, oriented generalized Heegaard surface obtained by thinning $H$. Such exists by Lemmas 4.2 and 4.4. By Corollary 5.10, $H$ is, up to isotopy, amalgamation-obtained from $\mathcal{H}^{\prime}$. Out of all generalized Heegaard surfaces obtained from $\mathcal{H}^{\prime}$ by a sequence of juggles and isotopies, let $\mathcal{H}$ be one which minimizes $|\mathcal{H}\cap\mathcal{S}|$. By Lemma 6.3, every coherent cycle in the dual digraph to $\mathcal{H}$ contains a thin sphere and, up to isotopy, $H$ is amalgamation-obtained from $\mathcal{H}$. We claim that $\mathcal{H}$ satisfies (3) and (4).
Claim 1: $\mathcal{S}$ is disjoint from $\mathcal{H}^{-}$
Suppose not. Consider an innermost curve $\zeta\subset\mathcal{S}\cap\mathcal{H}^{-}$ that bounds a disc $D\subset\mathcal{S}$ with interior disjoint from $\mathcal{H}^{-}$. Since $\mathcal{H}^{-}$ is incompressible, $D$ must be a semi-compressing disc for a component $F\subset\mathcal{H}^{-}$. Let $W\subset M\setminus\mathcal{H}^{-}$ be the component containing $D$ and let $J=\mathcal{H}^{+}\cap W$. As $\mathcal{H}$ is locally thin, by Lemma 3.6, $J$ can be isotoped so that each component of $J\cap D$ is inessential in $D$ and each innermost disc in $D$ is a semi-compressing disc for $H$.
Case 1a: $J\cap D=\varnothing$
In this case, by Lemma 3.3, the union of $D$ with a disc in $F$ is a sphere $Q$, which we can push slightly into the punctured 3-ball $B$ in $W$ bounded by $F\cup D$ and disjoint from $H$.
Let $A\subset M\setminus\mathcal{H}$ be the compressionbody on the opposite side of $F$ from $W$. Let $W^{\prime}\subset M\setminus\mathcal{H}^{-}$ be the component containing $A$. In $W^{\prime}$ we may perform an isotopy of the surface $H^{\prime}=\mathcal{H}^{+}\cap W^{\prime}$ so that it is still transverse to $\mathcal{S}$ and so that there is a path from $H^{\prime}$ to $Q$ which is disjoint from $\mathcal{S}$. (This isotopy may increase the number of intersections between $H^{\prime}$ and $\mathcal{S}$.) Using this path, perform a juggle of $F$ and $H^{\prime}$ over $Q$. After this juggle, we have not changed $|\mathcal{H}^{-}\cap\mathcal{S}|$, but the disc $D$ is now no longer a semi-compressing disc. We may therefore isotope $\mathcal{H}$ to eliminate it without increasing $|\mathcal{H}^{-}\cap\mathcal{S}|$. This contradicts our choice of $\mathcal{H}$.
Case 1b: $J\cap D\neq\varnothing$
Consider an innermost disc $E\subset D$ whose boundary is a component of $J\cap D$ that is innermost in $D$. As each component of $J\cap D$ is inessential in $J$, the union of $D$ with a disc $E\subset H$ is a sphere $Q$ bounding a punctured 3-ball in $M$ with interior disjoint from $J\setminus E$. Push $Q$ slightly into this 3-ball. As the interior of $D$ is disjoint from $\mathcal{H}^{-}$, there is an embedded path $\alpha$ with one endpoint on either $F$ or $J\setminus E$, one endpoint on $Q$, with interior intersecting $J$ in a single point contained in $E$, and which is otherwise disjoint from $\mathcal{H}\cup\mathcal{S}$. Let $R=J$. If $\alpha$ has an endpoint on $F$, let $P=F$; if it has an endpoint on $J$, let $P=J$. Use this path to perform a juggle of $R$ and $P$ over $Q$, as in Definition 6.1. After the juggle, $|\mathcal{H}\cap\mathcal{S}|$ is unchanged, but $D$ is no longer a semi-compressing disc. An isotopy of $J$ therefore removes it, decreasing $|J\cap\mathcal{S}|$. Repeating as necessary, we make $J$ disjoint from $D$. By Case 1a, we therefore contradict our choice of $\mathcal{H}$.
We conclude that $\mathcal{S}$ is disjoint from $\mathcal{H}^{-}$. The intersection of $\mathcal{S}$ with each component of $M\setminus\mathcal{H}^{-}$ is therefore the union of components of $\mathcal{S}$. Let $W\subset M\setminus\mathcal{H}^{-}$ be a component and let $J=\mathcal{H}^{+}\cap W$. By Lemma 3.6, we may isotope $J$ in $W$ so that for each component $S_{0}\subset\mathcal{S}$, either:
  • $S_{0}$ is a sphere and $J\cap S_{0}$ is the union of curves that are inessential in $J$ and the innermost such curves on $S_{0}$ bound semi-compressing discs for $J$.
  • $S_{0}$ is a compressing disc for $\partial M$ and the compressionbody of $W\setminus J$ containing $\partial S_{0}$ is a punctured product compressionbody. The intersection between $S_{0}$ and $J$ consists of curves that are inessential in $J$, the innermost ones on $S_{0}$ bound semi-compressing discs for $J$, and exactly one essential simple closed curve on $J$. That curve bounds a compressing disc for $J$ in the compressionbody on the opposite side of $J$ from that containing $\partial S_{0}$.
Suppose that some component of $\mathcal{S}\cap J$ is inessential in $J$. As in Case (1b) above, we may peform juggles, to eliminate those curves of intersection without increasing |±$|\mathcal{H}^{\pm}\cap\mathcal{S}|$. Consequently, we may assume that each component of $\mathcal{S}\cap W$ is either a sphere disjoint from $J$ or a compressing disc for $\partial M$ intersecting $J$ in a single simple closed curve, essential on $J$. This again contradicts our choice of $\mathcal{H}$, and so $\mathcal{H}$ must satisfy (3) and (4). ∎
7. The Strong Haken Theorem
We now prove the Strong Haken Theorem in the case when $\partial M$ does not contain spheres; the full statement follows in a straightforward fashion from this case.
Theorem 7.1 (Strong Haken).
Suppose that $H$ is a Heegaard surface for $M$, a compact, connected, orientable 3-manifold. Let $\mathcal{S}\subset M$ be a collection of pairwise disjoint essential spheres and discs. Then $H$ can be isotoped so that for each component $S_{0}\subset\mathcal{S}$, $H\cap S_{0}$ is a single simple closed curve.
The remainder of the paper is concerned with the proof of the Strong Haken Theorem. The bulk of the work lies in being careful how we amalgamate. For simplicity, assume that $\partial M$ does not have any spheres; the general case can be relatively easily obtained from this one and is covered in both [Scharlemann] and [HeSc]
Let $\mathbb{H}$ be the set of oriented generalized Heegaard surfaces for $M$ such that:
  1. (1)
    Each coherent cycle in $\mathcal{H}$ contains an edge corresponding to a thin sphere
  2. (2)
    $H$ is, up to isotopy, amalgamation-obtained from $\mathcal{H}$
  3. (3)
    Each sphere component of $\mathcal{S}$ is disjoint from $\mathcal{H}^{-}$ and is either disjoint from $\mathcal{H}^{+}$ or intersects $\mathcal{H}^{+}$ in a single simple closed curve that is essential on $\mathcal{H}^{+}$.
  4. (4)
    Each disc component of $\mathcal{S}$ is disjoint from $\mathcal{H}^{-}$ and intersects $\mathcal{H}^{+}$ in a single simple closed curve that is essential on $\mathcal{H}^{+}$.
  5. (5)
    Each compressionbody $C\subset M\setminus\mathcal{H}$ contains a collection of s-discs $\Delta_{C}$ such that after $\partial$-reducing $C$ along $\Delta_{C}$, each product compressionbody component contains a single scar and each component of $\mathcal{S}\cap\mathcal{H}$ is disjoint from that scar.
By Corollary 6.4, there is an $\mathcal{H}$ satisfying (1) - (3). By Lemma 3.4, we may choose the desired $\Delta_{C}$ so as to satisfy (4).
Without loss of generality, assume that out of all such $\mathcal{H}$ we choose $\mathcal{H}$ to have the fewest components. Note that this means that no component of $M\setminus\mathcal{H}$ is a product compressionbody bounded by a component of $\mathcal{H}^{+}$ and a component of $\mathcal{H}^{-}$. There may be such a product compressionbody incident to a component of $\partial M$. Let $P\subset\mathcal{H}^{-}$ be a minimal complete nonseparating collection of spheres. Let $f$ be a height function for $M\setminus P$.
Consider the sequence of amalgamations consistent with $f$ and spherical self-amalgamations that produce a Heegaard surface isotopic to $H$. If $\mathcal{H}\setminus P$ is connected, the sequence consists of only spherical self-amalgamations, otherwise it consists of some number of amalgamations followed by spherical self-amalgamations.
Begin by assuming that $\mathcal{H}\setminus P$ is disconnected. At each stage of the sequence, the generalized Heegaard surface satisfies (1) and (2). Consider the first step of this sequence and begin by supposing that it is an amalgamation of $J_{+}$ and $J_{-}$, each the union of thick surfaces, across $F$, the union of thin surfaces. Let $U_{\pm}$ be the (union of) compressionbodies on either side of $J_{+}$, with $U_{+}$ above $J_{+}$. Let $V_{\pm}$ be the (union of) compressionbodies on either side of $J_{-}$, with $V_{+}$ above $J_{-}$. Assume we have picked the notation so that $F=\partial_{-}U_{-}\cap\partial_{-}V_{+}$. If no component of $\mathcal{S}\cap U_{-}$ or $\mathcal{S}\cap V_{+}$ separates a component of $F$ from $J_{\pm}$, then the resulting generalized Heegaard surface still satisfies (1) - (5), and we contradict our choice of $\mathcal{H}$.
Consequently, some components of $\mathcal{S}$ separate $F$ from either $J_{-}$ or $J_{+}$. In this case, each such component of $\mathcal{S}$ is a sphere and the components of $F$ separated from $J_{-}$ and $J_{+}$ are also spheres. Suppose that $S_{1},\ldots,S_{k}\subset V_{+}$ separate components of $F$ from $J_{-}$. The exterior of $S_{1}\cup\cdots\cup S_{k}$ in $V_{-}$ is the disjoint union of compressionbodies, where one scar from each $S_{i}$ is a positive boundary component and one scar from each $S_{i}$ is a negative boundary component. Let $S^{\prime}_{1},\ldots,S^{\prime}_{k^{\prime}}\subset\mathcal{S}\cap U_{-}$ be those components separating components of $F$ from $J_{+}$. As before, their exterior is the union of compressionbodies such that each $S^{\prime}_{i}$ produces one scar in positive boundary and one in a negative boundary. Thus when we amalgamate $J_{-}$ and $J_{+}$ across $F$, using the discs $\Delta_{U_{-}}$ and $\Delta_{V_{+}}$, the resulting thick surface $J$ intersects each $S_{i}$ and $S^{\prime}_{i}$ exactly once in a curve parallel to a component of $\Delta_{U_{-}}$. If $J$ intersects each $S_{i}$ and $S^{\prime}_{i}$ exactly once, we do not need to perform further modifications. Otherwise, we do the following. For simplicity of exposition assume that each $S_{i}$ and $S^{\prime}_{i}$ intersects $J$ more than once. See Figure 8 for a depiction of the following isotopy.
Using the structure of the compressionbodies that are the complement of the union of the $S_{i}$ in $V_{+}$ and the $S^{\prime}_{i}$ in $U_{-}$, we may embed a collection of vertical discs $X$ in $(V_{+}|_{\Delta_{V_{+}}})\cup(U_{-}|_{\Delta_{U_{-}}})$ so that each disc is a square with one edge on $J_{-}$ or $J_{+}$, one edge on $\mathcal{S}$ and its other two edges on distinct tubes arising from the vertical extensions of the scars of $\Delta_{U_{+}}$. Furthermore, these discs may be constructed so that each intersects each $S_{i}$ and $S^{\prime}_{i}$ in at most 1 arc and so that the union of the arcs forms a tree with the components of $S_{i}\cap J$ and $S^{\prime}_{i}\cap J$ as the vertices. These discs guide an isotopy of $J$ so that after the isotopy it intersects each $S_{i}$ and each $S^{\prime}_{i}$ in a single simple closed curve. Since each $S_{i}$ and $S^{\prime}_{i}$ is essential, these simple closed curves are necessarily essential in $J$. As they are obtained by banding together the scars from $\Delta_{U_{-}}$, an isotopy ensures they are disjoint from the discs $\Delta_{U_{-}}\cup\Delta_{U_{+}}$. Afterwards, we have constructed a generalized Heegaard surface satisfying (1) - (5) but with fewer components than $\mathcal{H}$, a contradiction.
[…]
Figure 8. The solid black lines show the result of amalgamating two thick surfaces (shown as horizontal dashed and solid lines) across three thin spheres (shown in blue). The horizontal red lines denote two spheres in $\mathcal{S}$ and the discs $X$ are shown in green. Isotoping the amalgamated thick surface across $X$ ensures it meets each red sphere in a single simple closed curve.
Consequently, $\mathcal{H}$ is connected and we obtain $H$ (up to isotopy) by spherical self-amalgamations across a minimal set of nonseparating spheres $P$. When we perform the self-amalgamation, we first tube $\mathcal{H}$ to a parallel copy of $P$ and then tube $\mathcal{H}$ to itself via a tube that intersects each component of $P$ exactly once. Suppose that in $M\setminus P$, a component $S_{0}$ of $\mathcal{S}$ separated $\mathcal{H}$ from a scar of $P$. We may perform the initial tubing of $\mathcal{H}$ to $P$ so that the entirety of $S_{0}$ remains disjoint from $\mathcal{H}$ and then when we tube $\mathcal{H}$ to itself, each tube will intersect $S_{0}$ at most once. If some components of $\mathcal{S}$ intersect multiple tubes, an isotopy as in the previous paragraph ensures that we can arrange for $H$ to intersect each component of $\mathcal{S}$ in a single essential simple closed curve. ∎
Acknowledgements
The author is supported by a Colby College Research Grant and NSF Grant DMS-2104022.